# Human-Agent Joint Learning for Efficient Robot Manipulation Skill Acquisition

### v2

@@ Abstract
Employing a teleoperation system for gathering demonstrations offers the potential for more efficient learning of robot manipulation. However, teleoperating a robot arm equipped with a dexterous hand or gripper, via a teleoperation system poses significant challenges due to its high dimensionality, complex motions, and differences in physiological structure. In this study, we introduce a novel system for joint learning between human operators and robots, that enables human operators to share control of a robot end-effector with a learned assistive agent, facilitating simultaneous human demonstration collection and robot manipulation teaching. In this setup, as data accumulates, the assistive agent gradually learns. Consequently, less human effort and attention are required, enhancing the efficiency of the data collection process. It also allows the human operator to adjust the control ratio to achieve a trade-off between manual and automated control. We conducted experiments in both simulated environments and physical real-world settings. Through user studies and quantitative evaluations, it is evident that the proposed system could enhance data collection efficiency and reduce the need for human adaptation while ensuring the collected data is of sufficient quality for downstream tasks. Videos are available at https://norweig1an.github.io/human-agent-joint-learning.github.io/.
Keywords: Robotic Teleoperation, Robot Manipulation, Imitation Learning
footnotetext: * denotes equal contribution, †denotes corresponding author.
1 Introduction
The long-term vision in the field of robot learning has been to enable robots to perform diverse tasks at a human level in the physical world. Recently, significant progress has been made toward this goal by learning robot manipulation policies from demonstrations. Previous studies have utilized teleoperation systems [1, 2, 3, 4, 5, 6] to collect human demonstrations, and learning-based policies [7, 8, 9] have been formulated using the gathered data. Despite the notable advancements, several challenges still need to be addressed. For example, in vision-based teleoperation systems, even with state-of-the-art 3D hand pose estimation algorithms [10, 11, 12, 13], errors persist that significantly affect the teleoperation. Additionally, discrepancies between the structures of human hands and robot end-effectors, along with the lack of haptic feedback during contact-rich manipulation, also pose challenges. As a result, current teleoperation systems require human operators to practice extensively to adapt to these differences and gather the necessary data. This means that many human adaptations are essential. Furthermore, to meet the data requirements of robotic systems, humans need to collect large amounts of data, making this process very burdensome.
@@ 1 Introduction
Figure 1: Traditional frameworks typically separate human and agent training, requiring operators first to learn the task environment before data collection. This often leads to inefficiencies due to delayed and insufficient data gathering. In our framework, we integrate human and agent training from the start in a joint learning model. This enables simultaneous development and adapts the agents to human operation more effectively, enhancing overall efficiency and promoting better collaboration between humans and machines allowing for human effortless adaptation data collection.
@@
Naturally, a question was raised: in data-collection, how to make human adaptation less or even free while keeping the data quality? Here, we aim to address this question and argue that human-agent joint learning can help. That said, an effective and efficient teleoperation system should be designed to preferentially capture the operator’s intentions for directing a robot end effector and pose the main frame, while concurrently enabling an autonomous agent to help us ensure motion stability and interpolate the details. To this end, we propose a framework that achieves shared control between the human and a learned assistive agent. As shown in Fig. 1, our human effortless adaptation framework seeks to balance human needs and system performance, aiming to directly enhance the efficiency and quality of data collection by reducing the time and effort in human adaptation.
@@
Given our human-agent joint learning approach, we allow the data acquisition agent to grow and learn along with the human operator, reducing the learning burden of the human operator throughout the data collection task. Inspired by shared autonomy [14, 15, 16], we introduce a novel teleoperation system that enables collaboration between humans and learning-based agents to control a robot jointly. In particular, our proposed system provides the flexibility to adjust a “control ratio” between the human operator and a learning-based agent. A lower control ratio, in the beginning, signifies a greater emphasis on humans to teach the agent the finer-grained knowledge under the structure of human intention and principal actions. As the agent’s learning improves, a higher ratio later indicates increased autonomy from the learned agent to replace the human effort to “inpaint” the whole process given only human intention and principal actions. Additionally, once a sufficient amount of data is gathered, we have the option to transition the shared autonomy agent to full autonomy by reducing the human control ratio to zero.
We implement our framework as a joint learning system, which facilitates a human operator’s intention and an assistive agent’s execution of teleoperation tasks. For human operators, their inputs consist of intuitive control actions based on visual feedback and past cognitive experience. For the agent, their inputs are derived from a combination of sensors, data streams, and possibly pre-processed information. This can include visual data from cameras [1], tactile data from force sensors [3], and any additional context provided by the system [6], such as object recognition or environmental mapping. The agent processes these inputs using algorithms designed to interpret the task requirements and generate actions to assist the human operator, utilizing a diffusion model [8, 17, 18, 19] as the backbone enables us to adjust the control ratio by modifying the step number of the forward and reverse processes [20], providing a customized and adaptive approach to teleoperation tasks.
@@
We conducted experiments in six different simulation environments using two types of end-effectors: a dexterous hand and a gripper. Additionally, we performed experiments on three real-world tasks to validate our findings. Evaluation results indicate that our proposed system significantly enhances data collection efficiency, increasing the collection success rate by 30% and nearly doubling the collection speed. Additionally, data collected in shared autonomy mode is as effective for downstream tasks and models as data collected directly from the teleoperation system, demonstrating comparable validity. Our main contributions are summarized as follows:
@@
  • We study how to reduce human adaptation in teleoperation data collection and propose a human-agent joint learning paradigm.
@@
  • Conducting experiments to demonstrate the efficiency and effectiveness of our proposed system. Our system achieved significant performance improvements, including a $\mathit{30}\%$ increase in data collection success rate and nearly double the collection speed. We also deployed our system in a real-world environment and achieved significant results.
2 Related Works
@@
Teleoperation System. Data has always been a crucial foundation, and robots are no exception. Teleoperation serves as a significant source for collecting robot data [7, 21, 22, 23, 24, 25]. Some works achieve teleoperation through wearable devices [1, 2, 3, 4, 26], and vision-based teleoperation systems offer a low-cost and easily developed alternative [5, 6, 27, 28]. For instance, Li et al. [28] utilizes neural networks for markerless vision-based teleoperation of dexterous robotic hands from depth images. Handa et al. [5] set up a vision-based teleoperation system to control the Allegro Hand, accomplishing various contact-rich manipulation tasks in the real world. Recently, Qin et al. [6] introduced AnyTeleop, a unified teleoperation system designed to accommodate various arms, hands, realities, and camera setups within a singular framework. In this paper, we introduce a joint learning paradigm to assist teleoperation by sharing control between the human operator and a learning-based agent, aiming to improve the efficiency of the teleoperation process.
@@
Human Robot Cooperation. Collecting fine-grained human demonstration data for robotic manipulation is an effective but labor-intensive and time-consuming way to enable robots to complete a wide range of tasks [29, 30]. Previous work uses shared autonomy to assist people with disability in performing tasks by arbitrating human inputs and robot actions [31]. Many of the shared autonomy algorithms aim to estimate human intents from a set of pre-defined goals [32, 33, 34, 35] or by mapping low-dimension control input to high-dimension robot actions [31, 36]. In this work, we introduce a system that enables shared control between the human and assistive agent to facilitate the process of data collection and robot learning.
3 Technical Approach
The proposed system enables human operators to control the robot using a teleoperation system to gather training data (Sec. 3.1). Subsequently, utilizing the collected data, we train an agent (Sec. 3.2) to establish shared control between the human operator and the learned agent, thereby enhancing the efficiency of the data collection process (Sec. 3.3). Similar to the concept of “bootstrapping”, as more data accumulates, our system raises the control ratio of the learned agent, thereby reducing the effort required from human operators. This, in turn, enables us to collect even more data and continue improving the system iteratively. Moreover, we offer the option to transition the shared control agent to full autonomy once sufficient data is acquired (Sec. 3.4).
3.1 Teleoperation System.
Our pipeline initially captures the raw sensory signal $\mathcal{I}$. Human hand pose $\mathcal{P}^{h}\in\mathbb{R}^{20\times 3}$ can be obtained from the captured signal using off-the-shelf 3D hand pose estimation [10, 11, 13]. The pose $\mathcal{P}^{h}$ consists of the positions of $20$ keypoints. Then, employing an inverse kinematic function $f_{\textit{IK}}$, we compute the action of the robot $a\in\mathbb{R}^{m}$: $a=f_{\textit{IK}}(\mathcal{P}^{h}_{t},\mathcal{P}^{h}_{t+1})$, where it is calculated upon the change in the hand pose. Given this teleoperation system, the human operator will move the hand to produce a sequence of hand poses $\{\mathcal{P}^{h}_{i}\}_{i=0}^{T}$ to teleoperate the robot with an action sequence $\{a_{i}\}_{i=0}^{T}$ to achieve the task $\mathcal{T}$. The trajectory $\{(s_{i},a_{i})\}_{i=0}^{T}$ is the collected human demonstration data, where $s\in\mathbb{R}^{n}$ (here $n=18$) is the robot state, could be used for downstream tasks.
3.2 Diffusion-Model-Based Assistive Agent.
@@ 3 Technical Approach
After getting the data, we train a diffusion-model-based assistive agent to learn how to assist the human in collecting data in a shared control manner. We follow the Denoising Diffusion Probabilistic Model (DDPM) [18] training paradigm to construct the diffusion-model-based assist agent. The forward process of the Diffusion Model can be regarded as adding Gaussian noise to the data $x^{0}$ according to a variance schedule $\beta_{1:K}$ by
@@
where $p_{\theta}(x^{k-1}|x^{k})=\mathcal{N}(\mu_{\theta}(x^{k},k),\Sigma(x^{k},k))$ is usually referred as reverse process. Luo [38] shows that $p_{\theta}(x^{t-1}|x^{k})$ becomes tractable when conditioned on $x_{0}$ and Eq. 2 can be reformulated as minimizing the error in the noise prediction. Ho et al. [18] simplify the training loss function as
@@
Note that during the experiment, we found that adding slight noise to $s_{i}$ during training will give a better result. We assume doing so will augment the dataset.
3.3 Data Collection with Shared Control
Figure 2: To achieve shared control between the human and agent, we blend the action from the human operator $a^{h}$ using the forward and reverse process. The parameter $\gamma$ governs the control ratio, where a lower $\gamma$ results in the action better aligning with the human operator’s intention. In contrast, a higher $\gamma$ allows the learned agent to exert more influence over the blended action.
During data collection, the proposed system offers the option to control the robot in a shared control mode rather than directly applying the collected action $a^{h}$ from the teleoperation system. The classical shared autonomy method is achieved through the equation [32]:
@@ 3.3 Data Collection with Shared Control
where $a^{r}$ is generated by the learned agent. Considering that the agent operates as a diffusion policy (Fig. 2), we blend the action from the human with the forward and reverse processes. Given action $a^{h}$, a forward process diffuses the action as follows: $a^{k}=a^{h}+\epsilon^{k}.$ Subsequently, a reverse process denoises the action $a^{k}$:
@@
By applying action $a$, the control of the robot is shared between the human and the diffusion-model-based assistive agent. We can adjust the control ratio $\gamma=k/K$ between the human operator and the diffusion-model-based assistive agent by varying $k$. When $\gamma=0$, the action $a^{s}$ represents the teleoperation action $a^{h}$, which is the dexterous robot directly controlled by a human operator. As $\gamma$ approaches $1.0$, the action $a^{s}$ transitions to full autonomy $a^{r}$. A higher $\gamma$ value indicates a higher level of autonomy, allowing the learning-based agent more control rights to stabilize and direct the dexterous hand. This leads to a reduced human workload during the data collection process.
3.4 Intergrating Data Collection and Manipulation Learning.
We outline the overall process in Algo. 1. The assistive agent is trained in four steps as follows:
Step 2. Given the initial dataset, we train a less capable assistive agent to aid in further data collection. The training process has been formulated in Eq. 5 and Eq. 6, where a neural network $\epsilon_{\theta}$ is trained to predict noise $\epsilon$ out of the noisy action $a^{k}$.
Step 3. The trained agent assists in a second data collection round, aiming for higher efficiency and success. We refine the agent using data from both rounds to enhance its performance. This cycle repeats until the agent achieves full autonomy and the required data volume is collected.
@@
2:The collected dataset $\mathcal{D}$; the learned agent $f$; control ratio $\gamma$;
@@
5:     $\mathcal{H}$ collects data $d$ under $f$’s help; $\triangleright$ with proper control ratio $\gamma$, details shown in Fig. 2
@@
8:     Finetune $f$ with $\mathcal{D}$;
9:     Raise the agent’s control ratio $\gamma$;
10:return $\mathcal{D}$ and $f$;
Currently, the adjustment of the control ratio $\gamma$ is guided by intuitive assessments and careful monitoring of success rates. When an agent, trained with recently gathered data, exhibits a marked improvement in success rate relative to its prior performance, this enhancement is taken as a cue to fine-tune $\gamma$, thus refining control effectiveness. The assessment of the agent’s performance, followed by adjustments to $\gamma$, relies on empirical data and the agent’s performance in practice. This method ensures that decisions regarding the adjustment of $\gamma$ are firmly rooted in data, thus harmonizing the control strategy with the observed outcomes. It is posited that an agent reaches full autonomy when it achieves a success rate that aligns with or surpasses current state-of-the-art benchmarks.
4 Experiments
Tasks. We adopt six multi-stage manipulation tasks (Fig. 8). Pick-and-Place aims at picking an object on the table and placing it into a container. Articulated-Manipulation’s objective for the dexterous hand is to grasp and unscrew a door handle to open it, while for the gripper, it is to grab a drawer handle and pull the drawer open. Push-cube requires the robot to push the cube to the target position. Tool-Use aims at picking a hammer and using it to drive a nail into a board.
Efficiency of Data Collection. Our proposed system leverages shared control between human operators and learned agents to enhance the efficiency of data collection. To learn how the assistant agent could improve the data collection process, we conducted a user study.
Figure 8: Overview of six task settings and their task flow for Pick-and-Place (left), Articulated-Manipulation (middle), Gripper-Push (upper-right) and Dexterous-Tool-Use (bottom-right).
In the user study, 10 human operators participate, collecting data under two modes: one where control is shared between the operator and the learned agent (w/ Ours), and the other where control is directly by the operator alone (w/o Ours). Each participant is instructed to collect as much data as possible within 3 minutes under two different modes for 3 dexterous hand tasks. Three metrics are evaluated: Success Rate (Percent) indicates the percentage of attempts where data collection was successful. Horizon Length (Steps per Sample) measures the length of each collected trajectory, with a lower horizon length indicating smoother data collection. Collection Speed (Samples per Hour) refers to the number of successful trajectories that can be collected in one hour.

### v3

@@ Abstract
Employing a teleoperation system for gathering demonstrations offers the potential for more efficient learning of robot manipulation. However, teleoperating a robot arm equipped with a dexterous hand or gripper, via a teleoperation system presents inherent challenges due to the task’s high dimensionality, complexity of motion, and differences between physiological structures. In this study, we introduce a novel system for joint learning between human operators and robots, that enables human operators to share control of a robot end-effector with a learned assistive agent, simplifies the data collection process, and facilitates simultaneous human demonstration collection and robot manipulation training. As data accumulates, the assistive agent gradually learns. Consequently, less human effort and attention are required, enhancing the efficiency of the data collection process. It also allows the human operator to adjust the control ratio to achieve a trade-off between manual and automated control. We conducted experiments in both simulated environments and physical real-world settings. Through user studies and quantitative evaluations, it is evident that the proposed system could enhance data collection efficiency and reduce the need for human adaptation while ensuring the collected data is of sufficient quality for downstream tasks. For more details, please refer to our webpage https://norweig1an.github.io/HAJL.github.io/.
{strip}[Uncaptioned image]
@@
Figure 1: Human-agent joint learning overview. Traditional frameworks typically separate human and agent training, requiring operators first to learn the task environment before data collection. This often leads to inefficiencies due to delayed and insufficient data gathering. In our framework, we integrate human and agent training from the start in a joint learning model. This enables simultaneous development and adapts the agents to human operation more effectively, enhancing overall efficiency and promoting better collaboration between humans and machines allowing for human effortless adaptation data collection.
I INTRODUCTION
In recent years, significant progress has been made on learning robot manipulation policies from demonstrations. Previous studies have utilized teleoperation systems [1, 2, 3, 4, 5, 6] to collect human demonstrations, and learning-based policies [7, 8, 9] have been formulated using the gathered data. Despite the notable advancements, several challenges still need to be addressed. For example, in vision-based teleoperation systems, even with state-of-the-art 3D hand pose estimation algorithms [10, 11, 12, 13], errors persist that significantly degrade the teleoperation. Additionally, discrepancies between the structures of human hands and robot end-effectors, along with the lack of haptic feedback during contact-rich manipulation, also pose challenges. As a result, current teleoperation systems demand substantial human effort, in addition, collecting high-quality datasets remains a challenging and labor-intensive task in many scenarios.
@@ I INTRODUCTION
Naturally, a question was raised: in data-collection, how to make human effort less while improving the data quality? Here, we aim to address this question and argue that human-agent joint learning can help. That said, an effective and efficient teleoperation system should be designed to preferentially capture the operator’s intentions for directing a robot end effector and pose the main frame, while concurrently enabling an autonomous agent to help us ensure motion stability and interpolate the details. To this end, we propose a framework that achieves shared control between the human and a learned assistive agent during data collection. As shown in Fig. 1, our human-agent joint learning framework seeks to integrate the data collection and policy learning to enhance the efficiency of the whole process, reducing human effort, and improving the data quality.
@@
Given our human-agent joint learning approach, we allow the data acquisition agent to grow and learn along with the human operator. Inspired by shared autonomy [14, 15, 16], we introduce a novel teleoperation system that enables collaboration between humans and learning-based agents to control a robot jointly during the data collection and learning process. In particular, our proposed system provides the flexibility to adjust a “control ratio” between the human operator and a learning-based agent. A lower control ratio, in the beginning, emphasizes the human’s role in teaching the agent finer-grained knowledge under the structure of human intention and principal actions. As the agent’s learning improves, a higher ratio indicates greater autonomy from the learned agent to replace the human effort to “inpaint” the whole process given only human intention and principal actions.
@@
With the proposed system, the human effort will be reduced due to the shared control during data collection. Additionally, the agent learning process is integrated with the data collection, improving the efficiency of the whole process. In addition, the quality of the collected data is also improved, benefiting different kinds of downstream tasks. We conducted experiments in six different simulation environments using two types of end-effectors: a dexterous hand and a gripper. Additionally, we performed experiments on three real-world tasks to validate our findings. Evaluation results indicate that our proposed system significantly enhances data collection efficiency, increasing the collection success rate by 30% and nearly doubling the collection speed. Additionally, data collected in shared autonomy mode is as effective for downstream tasks and models as data collected directly from the teleoperation system, demonstrating comparable validity. Our main contributions are summarized as follows:
@@
  • We study how to reduce human adaptation while keeping data quality in teleoperation data collection and propose a human-agent joint learning paradigm.
@@
  • Conducting both simulation and real-world experiments to demonstrate the efficiency and effectiveness of our proposed system. Our system achieved significant performance improvements, including a 30% increase in data collection success rate and double the collection speed.
II RELATED WORKS
@@
Teleoperation for Data Collection. Data has always been a crucial foundation, and robots are no exception. Teleoperation serves as a significant source for collecting robot data [7, 17, 18, 19, 20, 21, 22]. Some works achieve teleoperation through wearable devices [1, 2, 3, 4, 23], and vision-based teleoperation systems offer a low-cost and easily developed alternative [5, 6, 24, 25]. For instance, [25] utilizes neural networks for markerless vision-based teleoperation of dexterous robotic hands from depth images. [5] set up a vision-based teleoperation system to control the Allegro Hand, accomplishing various contact-rich manipulation tasks in the real world. Recently, [6] introduced AnyTeleop, a unified teleoperation system designed to accommodate various arms, hands, realities, and camera setups within a singular framework. In this paper, we introduce a joint learning paradigm to assist teleoperation by sharing control between the human operator and a learning-based agent, improving the efficiency of data collection using teleoperation.
@@
Interactive robot learning. Collecting fine-grained human demonstration data for robotic manipulation is an effective but labor-intensive and time-consuming way to enable robots to complete a wide range of tasks [26, 27]. Previous work uses shared autonomy to assist people with disability in performing tasks by arbitrating human inputs and robot actions [28]. Many of the shared autonomy algorithms aim to estimate human intents from a set of pre-defined goals [29, 30, 31, 32], using clothoid curves to parametrize the state and control [33] or by mapping low-dimension control input to high-dimension robot actions [28, 34]. In this work, we introduce a system that integrates the agent’s learning process with data collection, facilitating both data collection and robot learning.
III PROPOSED METHOD
The primary contribution of this work is the development of a novel and highly efficient data collection method. To achieve this, the system is designed in two key stages: first, the proposed system allows human operators to control the robot via a teleoperation system to gather an initial but insufficient training dataset, which serves as the foundation for the second stage (Sec. III-B). Second, using these data, we train a diffusion-model-based assistive agent (Sec. III-C) to establish shared control between the human operator and the agent, thereby improving the efficiency of the data collection process (Sec. III-D). This approach mirrors the concept of “bootstrapping” [35], where, as more data is accumulated, the system progressively reduces the effort required from human operators, facilitating further data collection and iterative system improvement. Additionally, once sufficient data has been gathered, the system offers the option to transition the shared control agent to full autonomy.
III-A Preliminary
@@
To get a learned agent in Sec. III-C, enabling human-agent joint learning, we follow the Denoising Diffusion Probabilistic Model (DDPM) [36] training paradigm. Here we first briefly introduce the DDPM algorithm. The forward process of the Diffusion Model can be regarded as adding Gaussian noise to the data $x^{0}$ according to a variance schedule $\beta_{1:K}$ by
@@
where $p_{\theta}(x^{k-1}|x^{k})=\mathcal{N}(\mu_{\theta}(x^{k},k),\Sigma(x^{k},k))$ is usually referred as reverse process. [40] shows that $p_{\theta}(x^{t-1}|x^{k})$ becomes tractable when conditioned on $x_{0}$ and Eq. 2 can be reformulated as minimizing the error in the noise prediction. [36] simplify the training loss function as
@@
Figure 7: Simulation tasks overview. Here are six task settings and their task flow for Pick-and-Place (left), Articulated-Manipulation (middle), Gripper-Push (upper-right) and Dexterous-Tool-Use (bottom-right).
III-B Teleoperation System.
Our system initially captures the raw sensory signal $\mathcal{I}$. Human hand pose $\mathcal{P}^{h}$ can be obtained from the captured signal using off-the-shelf 3D hand pose estimation [10, 11, 13]. The pose $\mathcal{P}^{h}$ consists of the positions of the human hand’s keypoints. Then, employing an inverse kinematic function $f_{\textit{IK}}$, we compute the action of the robot $a\in\mathbb{R}^{m}$, such that $a=f_{\textit{IK}}(\mathcal{P}^{h}_{t},\mathcal{P}^{h}_{t+1})$, where it is calculated upon the change in the hand pose. Given this teleoperation system, the human operator will move the hand to produce a sequence of hand poses $\{\mathcal{P}^{h}_{i}\}_{i=0}^{T}$ to teleoperate the robot with an action sequence $\{a_{i}\}_{i=0}^{T}$ to achieve the task $\mathcal{T}$. The human collected demonstration trajectory $\{(s_{i},a_{i})\}_{i=0}^{T}$, where $s\in\mathbb{R}^{n}$ is the robot state, could be used for downstream tasks.
III-C Diffusion-Model-Based Assistive Agent.
After collecting data via teleoperation mentioned in Sec. III-B, we train a diffusion-model-based assistive agent to learn how to assist humans in collecting data in a shared control manner.
@@
Figure 8: Diffusion based shared control. To achieve shared control between the human and agent, we blend the action from the human operator $a^{h}$ using the forward and reverse process. The parameter $\gamma$ governs the control ratio, where a lower $\gamma$ results in the action better aligning with the human operator’s intention. In contrast, a higher $\gamma$ allows the learned agent to exert more influence over the blended action.
During data collection, the proposed system offers the option to control the robot in a shared control mode rather than directly applying the collected action $a^{h}$ from the teleoperation system. This leads to a reduced human workload during the data collection process. The classical shared autonomy method is achieved through the equation [29]:
@@
where $a^{r}$ is generated by the learned agent. However, considering that the agent operates as a diffusion policy (Fig. 8), we blend the action from the human with the forward and reverse processes. Given action $a^{h}$, a forward process diffuses the action as follows: $a^{k}=a^{h}+\epsilon^{k}.$ Subsequently, a reverse process denoises the action $a^{k}$:
@@
By applying action $a$, the control of the robot is shared between the human and the diffusion-model-based assistive agent. We can adjust the control ratio $\gamma=k/K$ between the human operator and the diffusion-model-based assistive agent by varying $k$. When $\gamma=0$, the action $a^{s}$ represents the teleoperation action $a^{h}$, which is the dexterous robot directly controlled by a human operator. As $\gamma$ approaches $1.0$, the action $a^{s}$ transitions to full autonomy $a^{r}$. A higher $\gamma$ value indicates a higher level of autonomy, allowing the learning-based agent more control rights to stabilize and direct the dexterous hand.
@@
2:The collected dataset $\mathcal{D}$; assistive agent $f$; control ratio $\gamma$;
@@
5:     $\mathcal{H}$ collects data $d$ under $f$’s help; $\triangleright$ see III-D2 for control ratio adjustment
@@
8:     end if
9:     Finetune $f$ with $\mathcal{D}$;
10:end while
11:return $\mathcal{D}$ and $f$;
III-D Integrating Data Collection and Manipulation Learning.
In this section, we show how to integrate data collection and manipulation learning into a unified framework that progressively reduces human effort and enhances robot autonomy.
III-D1 Detailed Algorithm Explanation
We outline the overall process in Algo. 1. The assistive agent is trained in three steps as follows:
Step 2. Given the initial dataset, we train a relatively low performance assistive agent to aid in further data collection. The training process has been formulated in Eq. 5 and Eq. 6, where a neural network $\epsilon_{\theta}$ is trained to predict noise $\epsilon$ out of the noisy action $a^{k}$.
Step 3. The trained agent assists in a second data collection round, aiming for higher efficiency and success. We then refine the agent using data from both rounds to enhance its performance. This cycle repeats until the agent achieves full autonomy and the required data volume is collected.
III-D2 Control Ratio Adjustment
For each data collection, we offer users two options to adjust the control ratio $\gamma$: (1) Users can empirically adjust $\gamma$ based on their needs. (2) Alternatively, set $\gamma=\frac{1}{2}(1+\cos\theta)$, where $\theta$ is obtained by calculating the dot product of the previous timestep’s human action $a^{h}$ and shared action $a^{s}$. This assesses alignment, increasing $\gamma$ for positive alignment to enhance agent control, and reducing it for misalignment to increase user control.
After obtaining the control ratio $\gamma$, we calculate the shared action $a^{s}$, using the human operator’s action $a^{h}$ as input, as defined in Eq. 8 (shown in Fig. 8).
IV EXPERIMENTS
In this section, we introduce the settings for both real world tasks and simulation tasks, along with the experimental results and data validation. Due to page limitations, we have included some detailed information such as training details and ablation study results on our webpage.
IV-A Tasks.
We adopt six multi-stage manipulation tasks (Fig. 7). Pick-and-Place aims at picking an object on the table and placing it into a container. Articulated-Manipulation’s objective for the dexterous hand is to grasp and unscrew a door handle to open it, while for the gripper, it is to grab a drawer handle and pull the drawer open. Push-cube requires the robot to push the cube to the target position. Tool-Use aims at picking a hammer and using it to drive a nail into a board.